# Survey on Question Answering over Visually Rich Documents: Methods, Challenges, and Trends

### v1

@@ Abstract
Using Large Language Models (LLMs) for Visually-rich Document Understanding (VrDU) has significantly improved performance on tasks requiring both comprehension and generation, such as question answering, albeit introducing new challenges. This survey explains how VrDU models enhanced by LLMs function, covering methods for integrating VrD features into LLMs and highlighting key challenges.
@@ 1 Introduction
Visually-rich documents (VRDs) combine complex information, blending text with visual elements like graphics, diagrams, and tables to convey detailed content effectively Ding et al. (2024a). Unlike traditional text documents, VRDs have two main features: text associated with typographic details (e.g., font, size, style, color), layout that organize information spatially, and visual elements, such as charts and figures, which enhance comprehension Huang et al. (2024a). These documents can be either native digital files (e.g., PDFs) containing searchable text and layout metadata, or scanned images requiring OCR to extract text and layout. Visually-rich Document Understanding (VrDU) is a rapidly evolving field at the intersection of computer vision and natural language processing, tackling both perception (document parsing, i.e. identification and extraction of objects within the document) and interpretation (analytical tasks on the document features, such as answering questions) Zhang et al. (2024c).
Current question-answering tasks on VrDU are not purely extractive; they require reasoning skills such as arithmetic, logical, and spatial reasoning, as well as counting, sorting, comparison, and multi-hop reasoning, which may involve combining multiple elements to answer a question, potentially spanning across modalities, pages, or even documents, as depicted in Table 1, illustrated by Figure 1.
Current approaches in document understanding have moved away from the traditional two-step pipeline of document parsing (converting the document into structured textual format) followed by interpretation (passing this structured text to a language model for answering questions) Huang et al. (2024b). Instead, they leverage a unified representation of the document that is directly processed by a LLM to tackle complex tasks. In this paradigm, the document parsing Zhang et al. (2024c) part becomes embedded within the model’s weights, with VrD encoders pretrained on document parsing tasks to learn representations that effectively capture all relevant document attributes Al-Shetairy et al. (2024), including contents (textual and visual), reading order, and specific layout information. We focus on this line of research in this paper.
To our knowledge, recent VrDU surveys focus on specific data types like charts Al-Shetairy et al. (2024) or tables and figures Huang et al. (2024b). The only surveys on whole documents (such as PDFs) are a catalog of models Ding et al. (2024b); Abdallah et al. (2024). More generally, no survey has thoroughly compared the different approaches to representing VrDs, specifically vision-only versus multi-modalities representations. So we provide a comprehensive analysis of how VrDU models interpret VrDs that often include multiple elements such as charts, tables, figures, and text, and span multiple pages (see Table 1), to answer questions. We break down the VrDU process into three key components: (1) Document representation – exploring encoding methods that incorporate text, layout, and visual data (Section 2) or vision-only approaches (Section 3), with a focus on multi-page (MP) representation (Section 4); (2) Embedding integration with LLMs – examining where and how document embeddings are incorporated (Section 5); and (3) Efficient training of these VrDU models, focusing on the pretraining approaches for modalities alignment (Section 6).
Datasets type #Pages #Tokens #Tab #Fig Crosspage Unans. Crossdoc #Regions Answer length
VisualMRC 2021 Wikipedia pages 1.0 151.46 ? ? 9.55
DocVQA 2021a Industry Documents 1.0 182.8 ? ? 2.43
InfographicVQA 2021b Posters (Canva, …) 1.2 217.9 ? ? 1.6
TAT-DQA 2022 Annual Companies Reports 1.3 550.3 >1 ? 3.44
MP-DocVQA 2023b Industry Documents 8.3 2026.6 ? ? 2.2
DUDE 2023 archives, wikimedia, … 5.7 1831.5 ? ? ✓(2.1%) ✓(12.7%) 3.4
MMLongBenchDoc 2024c ArXiv, Reports, Tutorials, … 47.5 21214.1 25.4% 20.7% ✓(33.0%) ✓(22.5%) 2.8
M3DocVQA 2024 Wikipedia pages 12.2 ? ? ? ? ✓(2.4k) ?
M-LongDoc 2024 ArXiv, Manuals, Reports 210.8 120988 71.8 161.1 180.3
MMDocBench 2024 Multi 1.0 ? ? ? 2.61 4.1
Table 1: Overview of Question-Answering VrDU datasets on PDFs or PowerPoint documents, summarizing Document Characteristics (e.g., average pages, tokens, tabs, figures per document) and Question Characteristics (e.g., presence of questions requiring cross-pages or cross-documents information, unanswerable questions, and average answer length). #Region refers to the number of regions identified for answering questions in datasets with coordinate annotations. Underlined datasets are standard benchmarks used for model comparison in Table 5.
@@ 2 Encoding VrDs from structured information
VrDs can be represented through three distinct but interconnected modalities: text and layout, derived from native digital formats or OCR extraction, and the overall visual appearance of the document, obtained by generating a screenshot of the document page. The most important layout features are bounding boxes around text and structural elements (e.g., tables). The visual modality captures the document page appearance, encompassing its overall structure and visual context of the document as a whole. The main problematic in VRD encoding is to represent and merge the information coming from these three distinct modalities. Table 2 summarizes encoder models VrD and that we describe in this section.
@@
ERNIE-Layout 2022 BERT Faster R-CNN emb. tables Disentangled attn (2021b) transformer
HiVT5 2022b T5 encoder DiT (2022) emb. tables VT5 encoder VT5 decoder
@@
RM-T5 2023 T5 encoder DiT (2022) emb. tables RMT (2022) T5 decoder
@@
Interaction of text and visual features with hierarchical / graph architectures
PICK 2020 transformer CNN GNN BiLSTM + CRF
GraphDoc 2022 SentenceBERT SwinT + FPN emb. tables GAT (2018) transformer
LAMPreT 2021 BERT CNN Hierarchical fusion transformer
@@ 2.1 Integrating the Layout information
The positions and sizes of elements within a document can vary in granularity, from individual tokens Łukasz Garncarek et al. (2021) to larger blocks like cells, tables, images, or paragraphs Li et al. (2021a, b). This layout information can be represented within VrDU models in three ways: through positional embeddings, attention bias, or directly within the text, as special tokens.
@@
The simplest approach, which does not require any architectural change, is to include layout information as special tokens, directly within the text. Lyrics Li et al. (2023c) and ViTLP Mao et al. (2024) add special tokens to represent layout information alongside text tokens. In Lyrics Li et al. (2023c), each word is tagged with its spatial coordinates using the $<Bbox>$ token, indicating its position within the document. ViTLP Mao et al. (2024) uses a $[LOC]$ token for the same purpose. The global text-layout sequence is based on an extended vocabulary $\hat{V}=V\cup[\text{LOC}]$, where $V$ is the original text vocabulary. This approach not only increases the sequence length, overloading the model’s context window, but also limits the ability to capture complex spatial interactions between elements in the document.
@@
This is why the VrDU community has focused on developing optimal methods to incorporate spatial information of tokens within documents. One way is to extend the 1D positional encoding of tokens in transformers to 2D (see Table 2) by embedding the spatial coordinates $(x,y)$ of each token’s bounding box. For example, LayoutLM Xu et al. (2019) embeds the $x$ and $y$ coordinates separately and sums them. DocFormer Appalaraju et al. (2021) further includes embeddings for the bounding box dimensions (height and width), while UNITER Chen et al. (2020) adds an embedding for the area of each bounding box. These embeddings can be learned or fixed (function-based, e.g., sinusoidal Hong et al. (2021)). Given the observation that 2D dependencies might vary across layers, some methods make the layout embeddings layer-specific Appalaraju et al. (2021). However, these approaches ignore the relative position of tokens within the document.
@@
To address this, some works extend the sequential positional attention bias used in seq-to-seq transformers, added before the softmax, to a 2D spatial context. This 2D positional bias represents the horizontal and vertical distances between tokens within the document Xu et al. (2022); Powalski et al. (2021). Fixed distance biases in attention are limited, as token semantics, like "total" in tables, often dictate specific spatial interactions beyond mere positional proximity. To ensure that the model pays particular attention to tokens located at the same horizontal position (as a token "total" for example), ERNIE-Layout Peng et al. (2022) introduces three relative position attention biases (disentangled attention), capturing respectively how the semantic meaning of a token interacts with its sequential, horizontal and vertical relative distance to the other token. FormNet Lee et al. (2022) goes further in this direction by allowing more complex interactions, using functions that combine semantic and position information, between the tokens semantics and relative positioning when computing the attention bias. This allows for instance to identify a label token positioned above or to the left of its corresponding value token.
For complex document layouts, methods that use positional biases in attention—especially semantic-based relative positional biases like the disentangled attention Peng et al. (2022) or the Rich Attention Lee et al. (2022)—seems to be more effective, at least on the Information Extraction tasks they have been compared on Jaume et al. (2019); Park et al. (2019).
@@ 2.2 Integrating the visual information
In all the reviewed works, the visual modality is a set of visual “tokens” (vectors), generated by a visual encoder. Initially based on CNNs Xu et al. (2022), these encoders have transitioned to Visual Transformers (ViTs) Huang et al. (2022).
@@
Fusing text and visual features for unified document encoding is challenging due to the differences between visual and text tokens (see Table 2).
Local modality alignment refers to aligning text and visual features specifically within localized regions of the document, focusing solely on the text and visual attributes from those regions. These regions can be either inferred using visual information, i.e. determined by an object detection module Carion et al. (2020); Ren et al. (2016) or determined by the textual information, i.e. considering the bounding boxes of text tokens Powalski et al. (2021). A simple method to locally align modalities involves summing the two representations per region Powalski et al. (2021). Note that regions without associated text only have a vision-only representation Tang et al. (2023).
Another line of works process differently visual and textual tokens representations. SelfDoc Li et al. (2021b) uses two cross-attentions for local alignment: from the visual to textual tokens and vice-versa, e.g. allowing the textual semantic representation to be contextualized by visual information such as color, bold elements, and position. For example, a large, bolded, centered text block is likely to serve as a title or header. By incorporating these visual cues, the model refines the semantic representation of text, ensuring that its meaning is informed by its visual context within the document. Rather than relying on costly cross-attention for modality fusion and interaction, Arctic-TILT Borchmann et al. (2024) introduces a lightweight attention mechanism after the transformer feed-forward layer to integrate visual information using a learnable role bias for text tokens, inspired by the TP-Attention Schlag et al. (2020).
@@
Global modality alignment involves considering both the visual and textual features of the entire document rather than on specific regions. A simple method to align those modalities globally involves concatenating them Xu et al. (2022). A transformer encoder then allows interaction through standard self-attention mechanisms Appalaraju et al. (2023); Huang et al. (2022). On the document scale, visual and textual tokens representations can also be processed differently. DocFormer Appalaraju et al. (2021) uses cross-attention for textual queries to guide the integration of visual information, and vice versa.
Finally, visual and textual tokens can interact by viewing the document as a graph, built thanks to the bounding boxes of elements like paragraphs, tables or figures, each representing a node. Graph-Based Relationship Modeling employs graph neural networks (GNNs) to capture relationships between modalities in documents Yu et al. (2020). Various graph-based approaches differ in how they aggregate neighbor information between nodes: Techniques include convolution (Graph Convolutional Networks) Yu et al. (2020), attention (Graph Attention Networks) Li et al. (2023d), and other aggregation methods Lee et al. (2022). While graph-based modeling is particularly effective at capturing long-range dependencies and contextual relationships in documents, it remains limited in facilitating interactions between the semantic of modalities, since modalities interaction is only layout-based.
@@ 3 Vision-Only Encoding of VrDs
DONUT Kim et al. (2021) 2560x1920 SwinT (2021a) MLP BART
@@
Pix2Struct Lee et al. (2023a) 1024x1024 ViT MLP BART
SeRum Lee et al. (2023b) 1280x960 SwinT 2021a MLP mBART
Kosmos2.5 Zheng et al. (2023) 224x224 Pix2Struct 2023a’s ViT Perceiver Resampler Transformer
@@
LLaVAR Zhu et al. (2023) 336x336 CLIP VIT-L/14 MLP Vicuna13B
Unidoc Nan et al. (2023) 336x336 CLIP VIT-L/14 MLP Vicuna13B
mPLUG-DocOwl Ye et al. (2023b) 224x224 CLIP VIT-L/14 Visual Abstractor Llama-7b
QwenVL Bai et al. (2023a) 448x448 CLIP-VIT-G/14 Cross-attn layer Qwen-7b
@@
SPHINX Chen et al. (2023c) 1344x896 VIT & ConvNext & MLP Llama2-7B
@@
UREADER Liu et al. (2023b) 2240x1792 CLIP ViT-L/14 MLP Vicuna13B
Monkey Yin et al. (2023) 1344x896 CLIP Vit-BigG Perceiver Resampler Qwen-7B
TextMonkey Yin et al. (2024) 1344x896 CLIP Vit-BigG Shared Perceiver Resampler Qwen-7B
mPLUG-DocOwl1.5 Li et al. (2024a) 2560x1920 EVA-CLIP H-Reducer Llama-7b + MAM
LLaVA-UHD Jiang et al. (2024) 672x1088 CLIP-ViT-L Shared perceiver Resampler Vicuna-13B
InternLMXC2-4KHD Hu et al. (2024b) 3840x1600 CLIP-ViT-L PLoRA matrix InternLM2-7B
Idefics2 Beurton-Aimar et al. (2024) 980x980 SigLIP-SO400M MLP Mistral-7B-v0.1
TextHawk Wang et al. (2024) 1344x1344 SigLIP-SO Perceiver Resampler InternLM-7B
TokenPacker Li et al. (2024c) 1344x1344 CLIP-ViT-L TockenPacker Vicuna-13B
mPLUG-DocOwl2 Zhang et al. (2024d) 504x504 EVA-CLIP H-Reducer+DocCompressor Llama-7b + MAM
@@
DocPedia Chen et al. (2023a) 2560x2560 SwinT 2021b MLP Vicuna-13B
LLaVA-PruMerge Liu et al. (2023a) 336x336 CLIP-ViT MLP Vicuna13B
CogAgent Chen et al. (2023b) 1120x1120 EVA2-CLIP & CogVLM Cross-attn layer & MLP Vicuna-13B
Vary Li et al. (2023a) 1024x1024 ViTDet & CLIP-ViT-L MLP Qwen-7B
Mini-Gemini Lu et al. (2024) 2048x2048 ConvNeXt & ViT-L/14 MLP Mistral-7B
LLaVA-HR Li et al. (2024b) 1024x1024 CLIP-ConvNeXt & ViT-L MLP & MR-Adapter Llama2-7B
TinyChart Xie et al. (2024a) 768x768 SigLIP MLP Phi-2
HRVDA not specified (2024) 1536x1536 SwinT (2021b) MLP Llama2-7B
@@
In the previous section, we discussed many different techniques to integrate visual and textual information. These models however remain complex pipelines because they need to deal with the different modalities.
Many recent works consider VrDs as images, which brings the advantage of dealing with only a single modality, relying on a (textual) decoder to handle different tasks. A summary of this type of model whose architecture we detail below is provided in Table 3.
Such approaches demand a highly capable visual encoder to capture all textual, layout, and visual details within the document. However, ViTs themselves are not capable to capture fine details within documents like text Anonymous (2024). Indeed, in ViTs, the visual input (e.g., a document page) is divided into fixed-size patches, each becoming a "vision token" (e.g., 14x14 or 16x16 pixels). If patches are too large, they may cover too much content, like multiple lines or text fragments, and miss fine details. Using smaller patches or increasing the image resolution creates more patches, enabling the model to capture finer details and better encode the document’s visual and textual content Lee et al. (2023a), but at the cost of efficiency.
Indeed, ViTs have a maximum context size (number of patches) they can manage Lee et al. (2023a). This is why research in vision-only VrDU focuses on architectural modifications to ViTs to enable the processing of high-resolution images (Section 3.1). A simpler alternative is to use a set of pre-trained ViTs, each handling a different part of the image, thereby allowing the processing of high-resolution images more efficiently (Section 3.2). In this case, the local of a global context might be harmful and needs to be tackled.

### v2

@@ Abstract
The field of visually-rich document understanding, which involves interacting with visually-rich documents (whether scanned or born-digital), is rapidly evolving and still lacks consensus on several key aspects of the processing pipeline. In this work, we provide a comprehensive overview of state-of-the-art approaches, emphasizing their strengths and limitations, pointing out the main challenges in the field, and proposing promising research directions.
Survey on Question Answering over Visually Rich Documents: Methods, Challenges, and Trends
Camille Barboule Orange Paris, France                        Benjamin Piwowarski Sorbonne Université, CNRS, ISIR Paris, France                        Yoan Chabot Orange Belfort, France
@@ 1 Introduction
Visually-rich documents (VRDs) combine complex information, blending text with visual elements like graphics, diagrams, and tables to convey detailed content effectively Ding et al. (2024). Unlike traditional text documents, VRDs have two main features: text associated with typographic details (e.g., font, size, style, color), layout that organize information spatially, and visual elements, such as charts and figures, which enhance comprehension Huang et al. (2024a). These documents can be either native digital files (e.g., PDFs) containing searchable text and layout metadata, or scanned images requiring OCR to extract text and layout. Visually-rich Document Understanding (VrDU) is a rapidly evolving field at the intersection of computer vision and natural language processing, tackling both perception (document parsing, i.e. identification and extraction of objects within the document) and interpretation (downstream tasks using the document features, such as answering questions or information extraction) Zhang et al. (2024c).
We provide a comprehensive analysis of how Visual Document Understanding (VrDU) models represent visually rich documents (VrDs) and use these features on downstream tasks, which often contain multiple elements—such as charts, tables, figures, and text—and span multiple pages (see Table 4 in appendix). Current VrDU approaches typically follow a two-step pipeline: document parsing followed by downstream tasks like question answering. We analyze how this two-step pipeline operates, looking first at how VrDU models encode VrDs, and then how large language models (LLMs) decode those features for downstream tasks.
We first take a deep dive into current approaches for processing and leveraging tokens and bounding boxes (extracted from OCR or PDF metadata) and linking textual and visual features within documents. Recent innovations aim to enable LLMs to handle the 2D positioning of elements in VrDs at different granularities and to process both textual and visual features from those documents, thereby improving their understanding of the structure and content of VrDs (Section 2).
Additionally, we examine how Large Vision-Language Models (LVLMs), which are increasingly recognized for their combined perception and reasoning capabilities, currently dominate the VrDU domain. Recent innovations focus on balancing coarse- and fine-grained visual representations of VrDs while limiting computational cost. Despite their growing popularity, we show that current LVLM architectures are still ill-suited to the specific challenges of VrDU, particularly in handling multi-page documents (Section 3).
Next, we analyze how VrDU approaches handle multi-page documents, exploring recent page-by-page strategies, strategies relying on sparse attention mechanisms to maintain connections across pages, and we finally examine retrieval-augmented generation (RAG) approaches that reduce the problem to a single-page context by retrieving relevant information from other pages, while giving insights on future promising directions (Section 4).
Finally, we compare the different approaches to optimally inject those visual information into a LLM to be processed optimally for downstream tasks, comparing self-attention and cross-attention-based approaches (Section 5).
@@ 2 Encoding VrDs from structured information
VrDs can be represented through three distinct but interconnected features: text and layout, derived from native digital formats or OCR extraction, and the overall visual appearance of the document, obtained by generating a screenshot of the document page. The most important layout features are bounding boxes around text and structural elements (e.g., tables). The visual modality captures the document page appearance, encompassing the overall structure and visual context of the document as a whole. The main problematic in VRD encoding is to represent and merge the information coming from these three distinct modalities. Table 1 summarizes models from this category that we detail in this section.
@@
ERNIE-Layout 2022 BERT Faster R-CNN emb. tables Disentangled attn (2021) transformer
HiVT5 2023 T5 encoder DiT (2022) emb. tables VT5 encoder VT5 decoder
@@
RM-T5 2024a T5 encoder DiT (2022) emb. tables RMT (2022) T5 decoder
@@ 2.1 Integrating the Layout information
The positions and sizes of elements within a document can vary in granularity, from individual tokens Garncarek et al. (2020); Xu et al. (2019) to larger blocks like cells, tables, images, or paragraphs Li et al. (2021a, b). This layout information can be represented within VrDU models in three ways: through absolute positional embeddings of the 2D position, as an attention bias / rotation depending on the spatial distance of the tokens, or directly within the text, as special tokens.
@@
The simplest approach, which does not require any architectural change, is to include layout information as special tokens, directly within the text Lu et al. (2024); Mao et al. (2024). The global text-layout sequence is based on an extended vocabulary $\hat{V}=V\cup[\text{BBOX}]$, where $V$ is the original text vocabulary. This approach not only increases the sequence length, overloading the model’s context window, but also limits the ability to capture complex spatial interactions between elements in the document.
@@
This is why the VrDU community has focused on developing optimal methods to incorporate spatial information of tokens within documents. One way is to extend the 1D absolute positional encoding of tokens in transformers to 2D (see Table 1) by embedding the spatial coordinates $(x,y)$ of each token’s bounding box. For example, LayoutLM Xu et al. (2019) embeds the discretized $x$ and $y$ coordinates separately and sums them. DocFormer Appalaraju et al. (2021) further includes embeddings for the bounding box dimensions (height and width), while UNITER Chen et al. (2020) adds an embedding for the area of each bounding box. These embeddings can be learned or fixed (function-based, e.g., sinusoidal Hong et al. (2022)).
@@
However, absolute positional encoding is limited, as they are added at the input only Chen et al. (2021). Recent models hence apply positional encoding directly within the attention mechanism for improved performance and flexibility. In particular, they extend the relative positional encoding Press et al. (2022); Raffel et al. (2023), applied on every self-attention layers, to a 2D space. Such approaches either encode the 2D distance as a bias term added before the softmax, representing the horizontal and vertical distances between tokens within the document Xu et al. (2022); Powalski et al. (2021), or as a rotation applied to the queries and keys vectors, depending on the absolute position of each token, inspired from 1D-RoPE Su et al. (2023), with a rotation of the attention score depending on the horizontal position of the token (e.g. position within a table row), and another on the vertical one (e.g. position within the columns of the table), with both scores weighted by a gating model Li et al. (2024a). Pondering the attention score with the 2D distance of the tokens is still limited, as token semantics, like "total" in tables, often dictate specific spatial interactions beyond mere positional proximity. To ensure that the model pays particular attention to tokens located at the same horizontal position of some meaningful tokens (like "total" in a table), ERNIE-Layout Peng et al. (2022) introduces three relative position attention biases (disentangled attention), capturing respectively how the semantic meaning of a token interacts with its sequential, horizontal and vertical relative distance to the other token. FormNet Lee et al. (2022) goes further in this direction by allowing more complex interactions, using functions that combine semantic and position information between tokens.
To conclude, in a world where documents are increasingly digital-native, with direct access to text and bounding boxes, enabling LLMs to handle such structures is crucial. However, the community has mostly focused on adapting either 1D absolute positional encodings or relative 1D positional bias to the 2D space, while little attention has been given to extending RoPE to 2D—despite most current models relying on it.
To the best of our knowledge, only a few studies focus on the granularity of positional information, distinguishing between intra-region positions (e.g., the position of a cell within a table or a token within a paragraph) and page-level positions (e.g., the position of a token or a region within the entire page). Region-level models fail to capture cross-region and word-level interactions, while page-level models (with token-wise positions) suffer from excessive contextualization Li et al. (2021b). We suggest that combining these two levels of granularity could enhance performance Wang et al. (2022).
@@ 2.2 Integrating the visual information
In all the works we reviewed, the visual modality is transmitted as a set of visual “tokens” (vectors), computed by a visual encoder. Initially based on CNNs Xu et al. (2022), these encoders have transitioned to Visual Transformers (ViTs) Huang et al. (2022).
@@
Fusing text and visual features for unified document encoding is challenging due to the differences between visual and text tokens (see Table 1). The integration of the two modalities can be done locally (per regions or the document) or globally (within the whole document).
@@
Global modality alignment involves considering both the visual and textual features of the entire document rather than specific regions. A simple method to align those modalities globally involves concatenating them Xu et al. (2022). A transformer encoder then allows interaction through standard self-attention mechanisms Appalaraju et al. (2023); Huang et al. (2022). However, such approaches require intensive pretraining for features (visual and textual) alignment Huang et al. (2022), since these two feature types form a unit within the document, sometimes representing the same elements (e.g., an image of a piece of text versus the text itself).
Local modality alignment refers to aligning text and visual features specifically within localized regions of the document, focusing solely on the text and visual attributes from those regions. These regions can be either inferred using visual information, i.e. determined by an object detection module Carion et al. (2020); Ren et al. (2016) or determined by the textual information, i.e. considering the bounding boxes of text tokens Powalski et al. (2021). A simple method to locally align modalities involves summing the two representations per region Powalski et al. (2021). Note that regions without associated text only have a vision-only representation Tang et al. (2023). However, this approach constrains the interaction between visual and textual modalities, thereby limiting the comprehensive understanding of the document region Li et al. (2021b).
To capture interactions between textual and visual features from a region of the document, SelfDoc Li et al. (2021b) uses two cross-attentions: from the visual to textual tokens and vice-versa, e.g. allowing the textual semantic representation to be contextualized by visual information such as color, bold elements, and position. For example, a large, bolded, centered text block is likely to serve as a title or header. By incorporating these visual cues, the model refines the semantic representation of text, ensuring that its meaning is informed by its visual context within the document. Rather than relying on costly cross-attention for modality fusion and interaction, Arctic-TILT Borchmann et al. (2024) introduces a lightweight attention mechanism after the transformer feed-forward layer to integrate visual information using a learnable role bias for text tokens, inspired by TP-Attention Schlag et al. (2020).
To conclude, the effect of the visual features, at least in the way it is utilized in such models (i.e. enriching the textual features’ representation), appears small and may primarily introduce redundancy to the textual elements: as shown by Tang et al. (2023), adding visual features brings little to no improvement on datasets without images or visual components, and only marginally enhances performance on highly visual tasks like InfographicsVQA Mathew et al. (2021a).
@@ 3 Vision-Only Encoding of VrDs
DONUT Kim et al. (2022) 2560x1920 SwinT (2021) MLP BART
@@
Pix2Struct Lee et al. (2023) 1024x1024 ViT MLP BART
SeRum Cao et al. (2023) 1280x960 SwinT 2021 MLP mBART
Kosmos2.5 Lv et al. (2024) 224x224 Pix2Struct 2023’s ViT Perceiver Resampler Transformer
@@
LLaVAR Zhang et al. (2024d) 336x336 CLIP VIT-L/14 MLP Vicuna13B
Unidoc Feng et al. (2023) 336x336 CLIP VIT-L/14 MLP Vicuna13B
mPLUG-DocOwl Ye et al. (2023a) 224x224 CLIP VIT-L/14 Visual Abstractor Llama-7b
QwenVL Bai et al. (2023) 448x448 CLIP-VIT-G/14 Cross-attn layer Qwen-7b
@@
SPHINX Lin et al. (2023) 1344x896 VIT & ConvNext & MLP Llama2-7B
@@
UREADER Ye et al. (2023b) 2240x1792 CLIP ViT-L/14 MLP Vicuna13B
Monkey Li et al. (2024d) 1344x896 CLIP Vit-BigG Perceiver Resampler Qwen-7B
TextMonkey Liu et al. (2024b) 1344x896 CLIP Vit-BigG Shared Perceiver Resampler Qwen-7B
mPLUG-DocOwl1.5 Hu et al. (2024a) 2560x1920 EVA-CLIP H-Reducer Llama-7b + MAM
LLaVA-UHD Xu et al. (2024a) 672x1088 CLIP-ViT-L Shared perceiver Resampler Vicuna-13B
InternLMXC2-4KHD Dong et al. (2024b) 3840x1600 CLIP-ViT-L PLoRA matrix InternLM2-7B
Idefics2 Laurençon et al. (2024) 980x980 SigLIP-SO400M MLP Mistral-7B-v0.1
TextHawk Yu et al. (2024) 1344x1344 SigLIP-SO Perceiver Resampler InternLM-7B
TokenPacker Li et al. (2024b) 1344x1344 CLIP-ViT-L TokenPacker Vicuna-13B
mPLUG-DocOwl2 Hu et al. (2024b) 504x504 EVA-CLIP H-Reducer+DocCompressor Llama-7b + MAM
@@
DocPedia Feng et al. (2024) 2560x2560 SwinT 2022b MLP Vicuna-13B
LLaVA-PruMerge Shang et al. (2024) 336x336 CLIP-ViT MLP Vicuna13B
CogAgent Hong et al. (2024) 1120x1120 EVA2-CLIP & CogVLM Cross-attn layer & MLP Vicuna-13B
Vary Wei et al. (2023) 1024x1024 ViTDet & CLIP-ViT-L MLP Qwen-7B
Mini-Gemini Li et al. (2024c) 2048x2048 ConvNeXt & ViT-L/14 MLP Mistral-7B
LLaVA-HR Luo et al. (2024) 1024x1024 CLIP-ConvNeXt & ViT-L MLP & MR-Adapter Llama2-7B
TinyChart Zhang et al. (2024b) 768x768 SigLIP MLP Phi-2
HRVDA Liu et al. (2024a) 1536x1536 SwinT (2022b) MLP Llama2-7B
@@
In the previous section, we discussed techniques that integrate visual and textual information. These models however remain complex because the segmentation between modalities in a document is not straightforward and may introduce redundancy, lead to information loss and require pretraining for modalities alignment.
Many recent works consider VrDs as images, which brings the advantage of dealing with a single modality, relying on a LLM decoder to handle different tasks. A summary of this type of model we detail below is provided in Table 2.
Such approaches, commonly named Large Visual-Language Models (LVLMs), demand a highly capable visual encoder to capture all textual, layout, and visual details within the document. However, ViTs themselves are not capable to capture fine details like text Zhang et al. (2025). Indeed, in ViTs, the visual input (e.g., a document page) is divided into fixed-size patches, each becoming a "vision token" (e.g., 14x14 or 16x16 pixels). If patches are too large, they may cover too much content, like multiple lines or text fragments, and miss fine details. Using smaller patches or increasing the image resolution creates more patches, enabling the model to capture finer details and better encode the document’s textual content Lee et al. (2023), but at the cost of efficiency.
Indeed, ViTs have a maximum context size (number of patches) they can manage Lee et al. (2023). This is why research in vision-only VrDU focuses on architectural modifications to ViTs to enable the processing of high-resolution images (Section 3.1). An effective alternative is to use a set of pre-trained ViTs, each handling a different part of the image, thereby allowing the processing of high-resolution images more efficiently (Section 3.2). In this case, it is necessary to ensure coherence between the cropped regions of the page.